Can someone help me on this problem?
I need to solve the following word problem and need to show the equation you use for the solution.
There are 850 Douglas fir and ponderosa pine trees in a section of forest bought by Sawz Logging Co. The company paid an average of $300 for each Douglas fir and $225 for each ponderosa pine. If the company paid $217,500 for the trees, how many of each kind did the company buy?
Thanks, Ashley

Answer by venugopalramana(3286)

Can someone help me on this problem?
There are 850 Douglas fir ( D SAY ) and ponderosa pine ( P SAY ) trees in a section of forest bought by Sawz Logging Co.
D+P=850.........1
D=850-P.............2
The company paid an average of $300 for each Douglas fir and $225 for each ponderosa pine.
COST OF D TREES =330D
COST OF P TREES = 225P
If the company paid $217,500 for the trees, how many of each kind did the company buy?
TOTAL COST = 330D+225P=217500.........3...SUBSTITUTING FOR D FROM EQN.2,WE GET
330(850-P)+225P=217500
-105P=217500-330*850
P=600
D=850-600=250